# Stepwise Regression and Pre-trained Edge for Robust Stereo Matching

### v3

@@ IV-C Ablation Study
In this section, we explore the effectiveness and optimal configuration of each component of SR-Stereo. Additionally, we apply the proposed stepwise architecture to existing methods to demonstrate its superiority in estimation performance and generalization performance.
TABLE I: Ablation study of DSR. The baseline is IGEV-Stereo. All methods run 15 disparity updates during inference. The $num_{GRU}$ denotes the usage times of ConvGRU, while the $num_{SRU}$ denotes the usage times of the proposed stepwise regression unit. The final configuration of $m$ is underlined. Bold: Best.
Experiments Variations SceneFlow Middlebury-H ETH3D Params.(M)
$num_{GRU}$ $num_{SRU}$ $m$ $Loss_{\Delta d}$ EPE(px) $>3px$ $>2px(\%)$ $>1px(\%)$
Baseline 15 0 - - 0.72 3.65 8.44 4.49 12.60
Range of Disparity Clip 0 15 1 - 0.71 3.61 8.42 4.10 12.77
0 15 2 - 0.69 3.51 7.89 4.47 12.77
0 15 3 - 0.70 3.55 7.66 4.44 12.77
0 15 4 - 0.69 3.54 7.31 5.14 12.77
Supervision of Disparity Clip 0 15 2 $\surd$ 0.70 3.49 6.78 4.05 12.77
0 15 3 $\surd$ 0.69 3.49 7.44 4.15 12.77
Stepwise Regression Unit (SRU) 4 11 2 $\surd$ 0.69 3.49 7.50 4.61 12.77
4 11 3 $\surd$ 0.70 3.56 7.80 4.61 12.77
8 7 2 $\surd$ 0.69 3.49 7.77 4.10 12.77
8 7 3 $\surd$ 0.70 3.52 7.91 4.79 12.77
IV-C1 Disparity Stepwise Regression
We explore the optimal settings for DSR as well as its effectiveness. Table I shows the results of DSR in different configurations. In the majority of configurations, the incorporation of DSR significantly enhances the performance of the baseline model across different datasets.
Firstly, experiments conducted with varying ranges (m) of disparity clips demonstrate that the choice of disparity clip range impacts the generalization performance on different datasets. For datasets with a small range of disparity, smaller disparity clips are preferred, while datasets with a larger range of disparity benefit from larger disparity clips. Interestingly, the DSR method achieves consistent and stable performance across all datasets when $m$ = 2.
Secondly, our method achieves further performance improvement through the supervision of disparity clips. By incorporating supervision specifically on disparity segments, the model’s accuracy and generalization capability are enhanced.
Furthermore, we investigate the impact of the number of stepwise regression units employed in the architecture. Experimental results reveal that increasing the usage times of stepwise regression units leads to better generalization performance of the model.
IV-C2 Optimal Configuration of Disparity Clip-Balanced Weight
We explore the optimal configuration of Disparity Clip-Balanced Weight. As shown in Table II, the utilization of Disparity Clip-Balanced Weight significantly improves the performance on multiple datasets. We observe that when the disparity is split into multiple clips, the imbalance problem is shifted from the distribution of disparity between different domains to the distribution of disparity clips within the same domain. Therefore, our proposed DSR and Disparity Clip-Balanced Weight is an effective combination for achieving excellent cross-domain generalization performance.
@@ IV-C3 Number of Stepwise Regression Units
Similar to the iteration-based methods, our SR-Stereo can trade off efficiency and performance by adjusting the number of update units. As shown in Table III, SR-Stereo can achieve better performance for the same number of updates compared to the best iteration-based method, IGEV-Stereo. It also shows that the range constraint on the update units does not reduce the convergence speed of the disparity, but rather makes the updated disparity more accurate.
IV-D Extension of the stepwise regression architecture
To further demonstrate the superiority of the stepwise architecture for disparity estimation and cross-domain generalization, we retain the existing update unit ConvCRU and apply the proposed Regression Objective Segmentation and Disparity Clip-Balanced Weight to the iteration-based methods RAFT-Stereo and IGEV-Stereo.
TABLE IV: Quantitative evaluation on SceneFlow and KITTI. Our SR-Stereo run 32 disparity updates during inference. Bold: Best.
Method SceneFlow KITTI 2015 KITTI 2012 Run-time(s)
D1-bg D1-fg D1-all 2-noc 2-all 3-noc 3-all EPE-noc EPE-all on KITTI
CREStereo [15] - 1.45 2.86 1.69 1.72 2.18 1.14 1.46 0.4 0.5 0.41
DLNR [22] 0.48 1.37 2.59 1.76 - - - - - - 0.28
Croco-Stereo [21] - 1.38 2.65 1.59 - - - - - - 0.93
UPFNet [27] - 1.38 2.85 1.62 1.67 2.17 1.09 1.45 0.4 0.5 0.25
PSMNet [12] 1.09 1.86 4.62 2.32 2.44 3.01 1.49 1.89 0.5 0.6 0.41
GANNet [41] 0.80 1.48 3.46 1.81 1.89 2.50 1.19 1.60 0.4 0.5 1.80
GwcNet [13] 0.98 1.74 3.93 2.11 2.16 2.71 1.32 1.70 0.5 0.5 0.32
AcfNet [42] 0.87 1.51 3.80 1.89 1.83 2.35 1.17 1.54 0.5 0.5 0.48
ACVNet [14] 0.48 1.37 3.07 1.65 1.83 2.35 1.13 1.47 0.4 0.5 0.20
RAFT-Stereo [16] 0.56 1.58 3.05 1.82 1.92 2.42 1.30 1.66 0.4 0.5 0.38
IGEV-Stereo [17] 0.47 1.38 2.67 1.59 1.71 2.17 1.12 1.44 0.4 0.4 0.18
SR-Stereo(ours) 0.45 1.37 2.49 1.56 1.66 2.07 1.09 1.36 0.4 0.4 0.19
Figure 7: Qualitative results on the test set of KITTI 2015. Both methods run 32 updates at inference. Our SR-Stereo is more accurate for edge regions and backgrounds.
TABLE V: Extension results for the stepwise regression architecture. For inference, IGEV-Stereo runs 15 disparity updates, while RAFT-Stereo runs 32 disparity updates. ROS:Regression Objective Segmentation. DCB: Disparity Clip-Balanced Weight ($h$=0.5). Gray: performance is improved after using the proposed method. Bold: Best.
@@ IV-D Extension of the stepwise regression architecture
As shown in Table V, the proposed method can improve the performance of existing iteration-based methods without changing the network structure. In addition, we can adjust the disparity clip range $m$ to achieve better results according to the characteristics of the existing methods. For instance, in the case of RAFT-Stereo, where the initial disparity is set to 0, a large clip range is required to accelerate the disparity convergence. Therefore, the best results are obtained with a clip range of $m$=5. Conversely, IGEV-Stereo obtains a coarse initial disparity through 3D-CNN, making a smaller clip range more suitable to achieve stable performance improvement.
IV-E Comparison with SOTA Methods
IV-E1 Benchmark Results
@@
In this section, we compare SR-Stereo with the state-of-the-art methods published on SceneFlow and KITTI. As described in Section IV-B1, we train the final version of SR-Stereo on SceneFlow. For the KITTI model, we fine-tune the final SceneFlow version of SR-Stereo for 50k steps. Tables IV shows the quantitative results. With similar training strategies, SR-Stereo achieves a new SOTA EPE on the SceneFlow test set. Evaluation results on the KITTI benchmark show that SR-Stereo achieves the best performance on the vast majority of metrics. Figure 7 shows a comparison of the qualitative results of SR-Stereo and IGEV-Stereo on KITTI 2015. Our method is more accurate for edge regions and backgrounds.
IV-E2 Zero-shot Generalization
We test the final SceneFlow version of SR-Stereo directly on Middlebury 2014 and ETH3D. As shown in Table VI, Our SR-Stereo achieves very competitive generalisation performance. Compared with the SOTA method IGEV-Stereo, our method achieves an overall improvement. The qualitative results are shown in Figure 8. As can be seen from the overall range of the disparity maps, our results are closer to the ground truth and show stronger generalization performance.
Figure 8: Generalization results on Middlebury 2014 and ETH3D. All methods run 32 disparity updates during inference. As can be seen from the overall range of the disparity maps, our results are closer to the ground truth and show stronger generalization performance.
@@ IV-E2 Zero-shot Generalization
TABLE VI: Synthetic data generalization experiments. The SR-Stereo run 32 disparity updates during inference. We pre-train our model on SceneFlow and test it directly on Middlebury 2014 and ETH3D. The 2-pixel error rate is used for Middlebury 2014, and 1-pixel error rate for ETH3D. Bold: Best.
Method Middlebury ETH3D
@@
PSMNet [12] 15.8 9.8 10.2
GANNet [41] 13.5 8.5 6.5
DSMNet [28] 13.8 8.1 6.2
STTR [7] 15.5 9.7 17.2
CFNet [26] 15.3 9.8 5.8
RAFT-Stereo [16] 8.7 7.3 3.2
IGEV-Stereo [17] 7.1 6.2 3.6
SR-Stereo(ours) 6.07 6.0 3.0
IV-F Domain Adaptation Based on Pre-trained Edge
As mentioned in Section III-A, in addition to proposing a robust network SR-Stereo, we also propose the DAPE to enhance the performance of existing models after being fine-tuned with sparse ground truth. In this section, we comprehensively evaluate the individual steps of DAPE and demonstrate its effectiveness through experimental results on KITTI, Middbury and ETH3D.
@@ IV-F Domain Adaptation Based on Pre-trained Edge
Figure 9: Edge Estimation Results for Different Inputs. From top to bottom: left RGB image, predicted edge map from the input of the left RGB image, disparity, predicted edge map from the concatenation of the left RGB image and disparity as a four-channel input, predicted Edge Map using our disparity estimator (see details in Figure 5). All methods are trained on the SceneFlow using the same network architecture (i.e., two residual blocks). Test samples are obtained from Middbury 2014 and KITTI-2015.
IV-F1 Different Inputs to the Edge Estimator
We explore the impact of different inputs on performance while utilizing the same edge estimator structure. Figure 9 shows the qualitative results that highlight the disparities between predicted edge maps obtained from different inputs. Notably, when the concatenation of disparity and RGB image is directly employed as input, the misleading influence of object surface texture is significantly reduced. This outcome demonstrates the effectiveness of integrating disparity information in edge estimation. Based on this, we further optimize the inputs of the disparity and RGB images to achieve accurate estimation of object edges using a lightweight network containing only two residual blocks.
@@ IV-F1 Different Inputs to the Edge Estimator
Figure 11: Visualization of ground truth sparse process for ETH3D and Middbury 2014. First, we predict the edge map using the edge estimator proposed in Section IV-F1. Then, we remove the pixels in the edge region of the ground truth and randomly remove the pixels in the non-edge region with a probability of 0.5.
IV-F2 Threshold for Edge Pseudo-label Generation
@@
As mentioned in Section III-C2, erroneous disparity estimation in ill-posed regions can lead to incorrect edge maps. We propose the use of threshold-based filtering to select pixels in non-edge region as pseudo-label. In this section, our focus is on exploring the threshold settings in different target domains. By doing so, we aim to analyze the impact of pseudo-label density on the adaptation of various domains. This analysis allows us to gain insights into the relationship between pseudo-label density and the effectiveness of DAPE.
@@
TABLE VII: Domain adaptation evaluation of DAPE on ETH3D. We experiment with the proposed DAPE on three models. All methods are fine-tuned on the sparsified ground truth and evaluated on the original ground truth. All methods run 15 disparity updates during inference. Gray: performance is improved after using DAPE. Bold: Best.
@@
We utilize sparsified ground truth for model fine-tuning and the original ground truth for evaluation, ensuring a more accurate assessment of the model’s domain-adaptive performance. The experimental results are presented in Table VII. As the threshold t decreases in DAPE, the density of generated edge pseudo-labels decreases as well. When $t$ is set to 0.25, all models exhibit performance improvement. This is due to the presence of reflective regions in some image pairs from ETH3D, which can lead to misleading edge map predictions. By using a smaller threshold, false edges are filtered out, resulting in more stable performance improvement. Figure 10 shows the qualitative results of DAPE for different models. It can be observed that our proposed DAPE effectively improves the performance of the model in the detail region.
TABLE VIII: Domain adaptation evaluation of DAPE on KITTI. We experiment with the proposed DAPE on three models. All methods run 15 disparity updates during inference. Gray: performance is improved after using DAPE. Bold: Best.
@@
TABLE IX: Domain adaptation evaluation of DAPE on Middbury 2014. All methods are fine-tuned on the sparsified ground truth and evaluated on the original ground truth. Due to the large disparity range of Middlebury 2014, RAFT-Stereo runs 32 disparity updates during inference, while the other two models run 15 disparity updates. We use the 2-pixel error as the evaluation metric. Gray: performance is improved after using DAPE. Bold: Best.
@@
As shown in Table VIII, the implementation of DAPE with most threshold values leads to performance improvements in both non-occluded and overall regions for all three models. Notably, the RAFT-Stereo model benefits significantly from higher-density edge pseudo-labels. Despite the presence of some erroneous labels in high-density edge pseudo-labels, they provide more comprehensive edge information, which effectively complements the edge disparity update guidance in RAFT-Stereo, especially considering its initial disparity is set to zero. Figure 12 showcases the qualitative results of DAPE on different models. Alongside enhancing the model’s performance in detailed regions, the proposed DAPE successfully mitigates disparity anomalies in textureless sky areas of the images.
@@
To verify the performance of DAPE at different resolutions, we evaluate it on Middlebury 2014 with three different resolutions. The qualitative results of DAPE on Middlebury 2014 are presented in Table IX. Consistent with the conclusions from the previous two datasets, the implementation of DAPE with most threshold values leads to performance improvements of the three models. This outcome further confirms the generalizability and robustness of DAPE across different scenarios.

### v4

@@ IV-C Ablation Study
In this section, we explore the effectiveness and optimal configuration of each component of SR-Stereo.
IV-C1 Stepwise Regression Architecture
We explore the optimal settings for stepwise regression architecture as well as its effectiveness. Table I shows the results of stepwise regression architecture in different configurations. In the majority of configurations, the incorporation of stepwise regression architecture significantly enhances the performance of the baseline model across different datasets.
In lines (b) to (e), experiments conducted with varying ranges (m) of disparity clips demonstrate that the choice of disparity clip range impacts the performance on different datasets. For datasets with a small range of disparity, smaller disparity clips are preferred, while datasets with a larger range of disparity benefit from larger disparity clips. Interestingly, the stepwise regression architecture achieves consistent and stable performance across all datasets when $m$ = 2.
In lines (f) to (g), our method achieves further performance improvement through the supervision of disparity clips. By incorporating supervision specifically on disparity clips, the model’s accuracy and generalization capability are enhanced.
In lines (h) to (k), we investigate the impact of the number of stepwise regression units employed in the architecture. Experimental results reveal that increasing the usage times of stepwise regression units leads to better generalization performance of the model.
IV-C2 Configuration of Disparity Clip-Balanced Weight
We explore the effectiveness of Disparity Clip-Balanced Weight. As shown in Table II, the utilization of Disparity Clip-Balanced Weight significantly improves the performance on multiple datasets. As mentioned previously, when the disparity error is split into multiple clips, the imbalance problem is shifted from the distribution of disparity between different domains to the distribution of disparity clips within the same domain. Therefore, our proposed stepwise regression architecture and Disparity Clip-Balanced Weight is an effective combination for achieving excellent performance.
@@ IV-C3 Number of Stepwise Regression Units
In the practice of this paper, SR-Stereo remains essentially an iteration-based method. Therefore, our SR-Stereo can trade off efficiency and performance by adjusting the number of update units. As shown in Table III, SR-Stereo can achieve better performance for the same number of updates compared to the baseline, IGEV-Stereo. This result also shows that the range constraint on the update units does not reduce the convergence speed of the disparity, but rather makes the updated disparity more accurate.
IV-C4 The Simplified Stepwise Regression Architecture
Notably, the proposed stepwise regression architecture contains three key components: the Stepwise Regression Unit, Regression Objective Segment (ROS), and Disparity Clip-Balanced Weight (DCB). Except for the Stepwise Regression Unit, the other two do not involve changes in the network structure. To further demonstrate the plug-and-play of the architecture, we apply the proposed ROS and DCB to the iteration-based methods RAFT-Stereo and IGEV-Stereo.
TABLE IV: Extension results for the simplified stepwise regression architecture. For inference, IGEV-Stereo runs 15 disparity updates, while RAFT-Stereo runs 32 disparity updates. ROS:Regression Objective Segment. DCB: Disparity Clip-Balanced Weight ($h$=0.5). Gray: performance is improved after using the proposed method. Bold: Best.
@@ IV-C4 The Simplified Stepwise Regression Architecture
TABLE V: Quantitative evaluation on SceneFlow and KITTI. Our SR-IGEV run 16 disparity updates during inference. Bold: Best. Underline: Second. Blue: Third. *: the baseline. Bottom right corner: Percentage improvement compared to the baseline.
Methods Years SceneFlow KITTI 2015 KITTI 2012 KITTI
D1-bg D1-fg D1-all 2-noc 2-all 3-noc 3-all EPE-all Time(s)
PSMNet [5] 2018 1.09 1.86 4.62 2.32 2.44 3.01 1.49 1.89 0.6 0.41
GANet [23] 2019 0.80 1.48 3.46 1.81 1.89 2.50 1.19 1.60 0.5 1.80
GwcNet [6] 2019 0.98 1.74 3.93 2.11 2.16 2.71 1.32 1.70 0.5 0.32
AcfNet [24] 2020 0.87 1.51 3.80 1.89 1.83 2.35 1.17 1.54 0.5 0.48
RAFT-Stereo [9] 2021 0.56 1.58 3.05 1.82 1.92 2.42 1.30 1.66 0.5 0.38
CREStereo [8] 2022 - 1.45 2.86 1.69 1.72 2.18 1.14 1.46 0.5 0.41
ACVNet [7] 2022 0.48 1.37 3.07 1.65 1.83 2.35 1.13 1.47 0.5 0.20
DLNR [13] 2023 0.48 1.37 2.59 1.76 - - - - - 0.28
Croco-Stereo [12] 2023 - 1.38 2.65 1.59 - - - - - 0.93
UPFNet [14] 2023 - 1.38 2.85 1.62 1.67 2.17 1.09 1.45 0.5 0.25
IGEV-Stereo* [10] 2023 0.47 1.38 2.67 1.59 1.71 2.17 1.12 1.44 0.4 0.18
NMRF [34] 2024 0.45 1.28 3.13 1.59 1.59 2.07 1.01 1.35 - 0.09
ADL-GwcNet [35] 2024 0.62 1.42 3.01 1.68 1.65 2.17 1.05 1.42 - -
DKT-IGEV [36] 2024 - 1.46 3.05 1.72 - - 1.22 1.56 - -
MoCha-Stereo [33] 2024 0.41 1.36 2.43 1.53 - - 1.06 1.36 0.4 0.34
Selective-IGEV [11] 2024 0.44 1.33 2.61 1.55 1.59 2.05 1.07 1.38 0.4 0.24
[…]
Figure 10: Qualitative results on the test set of KITTI 2015. Both methods run 32 updates at inference. Our SR-IGEV is more accurate for edge regions and backgrounds.
@@
As shown in Table IV, the ROS and DCB can improve the performance of existing iteration-based methods without changing the network structure, and therefore can be generalized very easily. In addition, we can adjust the disparity clip range $m$ to achieve better results according to the characteristics of the existing methods. For instance, in the case of RAFT-Stereo, where the initial disparity is set to 0, a large clip range is required to accelerate the disparity convergence. Therefore, the best results are obtained with a clip range of $m$=5. Conversely, IGEV-Stereo obtains a coarse initial disparity through 3D-CNN, making a smaller clip range more suitable to achieve stable performance improvement.
IV-D Comparison with SOTA Methods
IV-D1 Benchmark Results
@@
In this section, we compare SR-Stereo with the state-of-the-art methods published on SceneFlow and KITTI. As described in Section IV-B1, we train the SR-IGEV on SceneFlow for 200k steps. Then, we fine-tune the SR-IGEV on the KITTI for 50k steps. Tables V shows the quantitative results. As shown in Table V, unlike the state-of-the-art generalized stereo method DKT-IGEV [36], our method does not deteriorate the in-domain performance of the baseline model. On the contrary, surprisingly, with a similar training strategy, our SR-IGEV achieves similar performance to the state-of-the-art method Selective-IGEV [11]. More importantly, compared to the baseline IGEV-Stereo, our method only increases the time cost by 0.01s, which is far better than the two IGEV-based SOTA methods (Selective-IGEV [11] and MoCha-Stereo [33]). All these results effectively demonstrate the advantages of our method. Figure 10 shows a comparison of the qualitative results of SR-Stereo and IGEV-Stereo on KITTI 2015. Our method is more accurate for edge regions and backgrounds.
Figure 11: Generalization results on Middlebury 2014 and ETH3D. All methods run 32 disparity updates during inference. Our SR-IGEV performs better in the background and detail regions.
IV-D2 Zero-shot Generalization
We test the SR-IGEV directly on Middlebury, ETH3D and KITTI. As shown in Table VI, Our method achieves very competitive generalization performance. Compared with the baseline IGEV-Stereo, our method achieves an overall improvement. Moreover, compared to the majority of generalized stereo methods, our method achieves the best performance on Middlebury and ETH3D. The generalization performance of our method on KITTI is also very competitive among the methods of the last two years. We also point out that we recommend evaluating the generalization performance more objectively based on the results on Middlebury and ETH3D, as is done in IGEV-Stereo [10] and MoCha-Stereo [33]. Because the ground truth in KITTI is too sparse, which ignores errors in some regions. Finally, we visualize the generalization results of SR-IGEV in Figure 11. It can be seen that our method performs better in the background and detail regions.
@@ IV-D2 Zero-shot Generalization
TABLE VI: Cross-domain generalization evaluation. Our SR-IGEV run 32 disparity updates during inference. We pre-train our model on SceneFlow and test it directly on Middlebury 2014, ETH3D and KITTI. The 2-pixel error rate is used for Middlebury 2014, 1-pixel error rate for ETH3D and D1-error for KITTI. Bold: Best. Underline: Second. *: the baseline. Bottom right corner: Percentage improvement compared to the baseline.
Methods Type Years Middlebury ETH3D KITTI-12 KITTI-15
@@
PSMNet [5] 2018 15.8 9.8 10.2 15.1 16.3
GANet [23] 2019 13.5 8.5 6.5 10.1 11.7
DSMNet [15] 2020 13.8 8.1 6.2 6.2 6.5
STTR [4] 2021 15.5 9.7 17.2 - -
RAFT-Stereo [9] Normal 2021 8.7 7.3 3.2 5.7 5.5
IGEV-Stereo* [10] 2023 7.1 6.2 3.6 5.2 6.0
MoCha-Stereo [33] 2024 6.2 4.9 3.2 - -
Selective-IGEV [11] 2024 6.8 6.6 5.4 5.6 6.0
NMRF [34] 2024 - 7.5 3.8 4.2 5.1
Graft-GANNet [37] 2022 - - 6.2 4.9 4.2
FC-DSMNet [38] 2022 12.0 7.8 6.0 5.5 6.2
HVT-GwcNet [39] Generalized 2023 10.3 - 5.9 3.9 5.0
Mask-LacGwcNet [40] 2023 16.9 - 5.3 5.7 5.6
ADL-GwcNet [35] 2024 9.1 - 3.8 4.5 4.2
DKT-IGEV [36] 2024 6.9 - 3.1 3.8 4.8
SR-IGEV(Ours) - - $\textbf{6.0}_{{\color[rgb]{1,0,0}-15.5\%}}$ $\underline{6.0}_{{\color[rgb]{1,0,0}-3.2\%}}$ $\textbf{3.0}_{{\color[rgb]{1,0,0}-16.7\%}}$ $4.9_{{\color[rgb]{1,0,0}-5.8\%}}$ $5.7_{{\color[rgb]{1,0,0}-5\%}}$
@@
Figure 12: Edge Estimation Results for Different Inputs. (a) Based on RGB image only. (b) Based on disparity only. (c) Based on concatenation of RGB image and disparity. (d) Step-by-step input of RGB image and disparity (Figure 8). All methods are trained on the SceneFlow using the same network architecture (i.e., two residual blocks in Figure 8). Test samples are obtained from Middbury 2014 and KITTI-2015.
IV-E Domain Adaptation based on Pre-trained Edge
As mentioned in Section III-A, in addition to proposing a robust network SR-Stereo, we also propose the DAPE to enhance the performance of existing models fine-tuned with sparse ground truth. In this section, we comprehensively evaluate the individual steps of DAPE and demonstrate its effectiveness through experimental results on KITTI, Middbury and ETH3D.
IV-E1 Different Inputs to the Edge Estimator
We explore the impact of different inputs on performance while utilizing the same edge estimator structure (Figure 8). Figure 12 shows the qualitative results that highlight the differences between predicted edge maps obtained from different inputs. When only RGB images are used, the edge maps are excessively noisy due to the interference of object surface textures. In contrast, with the introduction of disparity, the noise in the edge map is significantly reduced. To further demonstrate the impact of edge detection on DAPE, we fine-tune IGEV-Stereo using different edge results. As shown in Table VII, the edge maps based only on RGB images worsen the effect of the DAPE architecture, which is particularly obvious on ETH3D. In addition, the three disparity-based edge maps do not show a significant difference in the DAPE architecture, and all of them improve the fine-tuning performance. However, considering the visualization in Figure 12 and the diversity of signal sources, we recommend using the approach in Figure 8.
TABLE VII: DAPE fine-tuning results based on different edge maps. The baseline is IGEV-Stereo and is fine-tuned on the sparsified ground truth (see Figure 13 for more details). ED: Edge Detection. $\dagger$: Based on concatenation of RGB image and disparity. *: Step-by-step input of RGB image and disparity (Figure 8). The $t=0.25$ in DAPE. Bold: Best.
Methods ED Input Middlebury ETH3D
RGB Disparity Full Half
Baseline - - 11.72 5.14 1.65
DAPE $\surd$ - 11.76 5.01 2.05
- $\surd$ 11.69 4.89 1.62
[…]
$\surd$*√ ∗ $\surd$*√ ∗ 11.69 4.93 1.61
@@ IV-E1 Different Inputs to the Edge Estimator
Figure 13: Visualization of ground truth sparse process for ETH3D and Middbury 2014. First, we predict the edge map using the edge estimator proposed in Section IV-E1. Then, we remove the pixels in the edge region of the ground truth and randomly remove the pixels in the non-edge region with a probability of 0.5.
IV-E2 Threshold for Edge Pseudo-label Generation
@@
As mentioned in Section III-D2, erroneous disparity estimation in ill-posed regions can lead to incorrect edge maps. We propose the use of threshold-based filtering to select pixels in non-edge region as pseudo-label. In this section, our focus is on exploring the threshold settings in different target domains. By doing so, we aim to analyze the impact of pseudo-label density on the adaptation of various domains. This analysis allows us to gain insights into the relationship between pseudo-label density and the effectiveness of DAPE.
@@
TABLE VIII: Domain adaptation evaluation of DAPE on ETH3D. We experiment with the proposed DAPE on three models. All methods are fine-tuned on the sparsified ground truth and evaluated on the original ground truth. All methods run 15 disparity updates during inference. Gray: performance is improved after using DAPE. Bold: Best.
@@
We utilize sparsified ground truth for model fine-tuning and the original ground truth for evaluation, ensuring a more accurate assessment of the model’s domain-adaptive performance. The experimental results are presented in Table VIII. As the threshold t decreases in DAPE, the density of generated edge pseudo-labels decreases as well. When $t$ is set to 0.25, all models exhibit performance improvement. This is due to the presence of reflective regions in some image pairs from ETH3D, which can lead to misleading edge map predictions. By using a smaller threshold, false edges are filtered out, resulting in more stable performance improvement. Figure 14 shows the qualitative results of DAPE for different models. It can be observed that our proposed DAPE effectively improves the performance of the model in the detail region.
TABLE IX: Domain adaptation evaluation of DAPE on KITTI. We experiment with the proposed DAPE on three models. All methods run 15 disparity updates during inference. Gray: performance is improved after using DAPE. Bold: Best.
@@
TABLE X: Domain adaptation evaluation of DAPE on Middbury 2014. All methods are fine-tuned on the sparsified ground truth and evaluated on the original ground truth. Due to the large disparity range of Middlebury 2014, RAFT-Stereo runs 32 disparity updates during inference, while the other two models run 15 disparity updates. We use the 2-pixel error as the evaluation metric. Gray: performance is improved after using DAPE. Bold: Best.
@@
As shown in Table IX, the implementation of DAPE with most threshold values leads to performance improvements in both non-occluded and overall regions for all three models. Notably, the RAFT-Stereo model benefits significantly from higher-density edge pseudo-labels. Despite the presence of some erroneous labels in high-density edge pseudo-labels, they provide more comprehensive edge information, which effectively complements the edge disparity update guidance in RAFT-Stereo, especially considering its initial disparity is set to zero. Figure 15 showcases the qualitative results of DAPE on different models. Alongside enhancing the model’s performance in detailed regions, the proposed DAPE successfully mitigates disparity anomalies in textureless sky areas of the images.
@@
To verify the performance of DAPE at different resolutions, we evaluate it on Middlebury 2014 with three different resolutions. The qualitative results of DAPE on Middlebury 2014 are presented in Table X. Consistent with the conclusions from the previous two datasets, the implementation of DAPE with most threshold values leads to performance improvements of the three models. This outcome further confirms the generalizability and robustness of DAPE across different scenarios.